10. Give possible upressions for the length and breadth of each of the folwwing rectangles, in which their areas are given:
I Area: 26o! - a5a • 12 I I Area: asr. • 1ay - 12 I
(i) (ii")

Grade:12th pass

1 Answers

Pawan Prajapati
3 years ago
Area = 2002 - 35a + 12 = 25a2 - 20a - 15a + 12 = 5a(5a - 4) - acoo - 4) = coo - 3)(5a - 4). Possible expressions for length and breadth are (5a - 3) and (5a - 4) respectively. (ii) Area = 35y2 + 13y - 12 = 35y2 + 28y - 15y - 12 = 7y(5y + 4) - 3(6y + 4) = (7y - 3)(5y + 4). Possible expressions for length and breadth are (7y - 3) and (5y + 4).